# Simple Guidance Mechanisms for Discrete Diffusion Models

### v1

@@ 5.2 Guided Discrete Diffusion
AR FUDGEγ=7 924.4 $\pm$ 12.05 53.0 $\pm$ 3.54 0.61 $\pm$ 0.00
AR PPLMn=30,η=0.1 1007.2 $\pm$ 2.28 142.0 $\pm$ 14.2 0.45 $\pm$ 0.00
[…]
Figure 3: Diffusion models extend the steer-ability Pareto frontier compared to AR. (Left) Comparing D-CBG and FUDGE for maximizing the drug-likeness (QED) property in the QM9 dataset. Using our D-CBG outperforms FUDGE classifier guidance for AR. Only UDLM can accommodate larger $\gamma$ values that enable better molecule property maximization. (Right) Comparing D-CFG for diffusion and AR models for maximizing ring-count in the QM9 dataset. Diffusion models better trade-off novel generation and property maximization.
[…]
Method Guidance Valid ($\uparrow$) Novel ($\uparrow$) Ring Count Mean ($\uparrow$)
Original Data - 133k 133k 1.74
Classifier-free
AR D-CFGγ=5 441.4 $\pm$ 11.08 77.8 $\pm$ 2.28 4.83 $\pm$ 0.08
MDLM D-CFGγ=5 90.0 $\pm$ 8.15 60.0 $\pm$ 7.52 5.26 $\pm$ 0.15
UDLM D-CFGγ=3 998.2 $\pm$ 4.49 216.2 $\pm$ 12.99 4.88 $\pm$ 0.04
Classifier-based
AR FUDGEγ=5 281.2 $\pm$ 4.97 103.6 $\pm$ 4.83 4.89 $\pm$ 0.1
AR PPLMn=10,η=0.1 1009.8 $\pm$ 1.92 140.6 $\pm$ 16.47 1.92 $\pm$ 0.09
[…]
@@
Table 7: Guidance improves quality on discretized CIFAR10. FID and IS for finite- (D3PM) and continuous-time (MDLM / UDLM) discrete diffusion models. Guidance using D-CFG ($\gamma=4$). Best values are bolded. Taken from Austin et al. (2021).
FID ($\downarrow$) IS ($\uparrow$) D3PM Absorb 41.28 6.26 MDLM 33.75 6.74 MDLM D-CFG 15.56 9.02 D3PM Uniform 51.27 5.99 UDLM 33.65 6.86 UDLM D-CFG 23.21 8.66
Figure 4: CIFAR10 conditional generation.
Ablation: Faster Sampling  In Table 8, we explore using faster inference settings (smaller $T$) where diffusion models predict multiple pixels in parallel. These results focus on the CIFAR10 dataset; see Appendix D.3 for additional experiments on other datasets. MDLM’s performance deteriorates with smaller $T$, whereas UDLM is robust to this setting. This validates a key motivation behind UDLM: in settings where MDLM ‘locks in’ certain predictions that it cannot change, UDLM is more resilient given that all tokens can change throughout the decoding process.
@@ 6 Related Works, Discussion, and Conclusion
Discrete Diffusion  Recent works have examined interpolating discrete diffusion (Ou et al., 2024; Sahoo et al., 2024a; Shi et al., 2024; Zhao et al., 2024; Zheng et al., 2023), a special case of the general framework proposed in D3PM (Austin et al., 2021). Our UDLM method most closely aligns to the recent state-of-the-art discrete diffusion models of Ou et al. (2024), Sahoo et al. (2024a), and Shi et al. (2024) that focus on absorbing state diffusion. Similar to these works, we provide a continuous time ELBO leading to performance gains, but we focus on uniform noise. Of note, other extensions of D3PM that do not start from the variational perspective, instead rely on the formalisms of continuous time Markov chains (CTMC) (Campbell et al., 2022) and concrete score matching (Lou et al., 2023), but are less performant than works such as MDLM (Sahoo et al., 2024a) and our method. Also stemming from CTMC are extensions of flow-based (Lipman et al., 2022) approaches to discrete data (Campbell et al., 2024; Gat et al., 2024). In Appendix B, we discuss connections of our work to CTMC in more detail.
Guidance  Leveraging continuous embeddings of discrete data, Diffusion-LM (Li et al., 2022) uses Langevin sampling with classifier-based guidance. Similarly, SSD-LM (Han et al., 2022) perform Gaussian noising on the logits of a bi-directional model, which they combine with pre-trained classifiers to perform classifier-based guidance with Langevin dynamics. LD4LG (Lovelace et al., 2024) implement classifier-free guidance on continuous embeddings. Stark et al. (2024) use flow-matching on the simplex to perform classifier-free and classifier-based guidance on continuous representations. Wang et al. (2023) perform guidance using auxiliary semantic latent variables. In contrast, to these continuous formulations for discrete data, our work adapts guidance mechanisms directly to the discrete domain.
Table 8: UDLM is robust to faster sampling. Images sampled from a conditional model (D-CFGγ=1) trained on CIFAR10. F1 metric from a separate classifier trained to identify class label used for generation. Best metric per $T$ is bolded.
Model FID ($\downarrow$) IS ($\uparrow$) F1 ($\uparrow$)
@@
MDLM 27.94 7.14 0.81
UDLM 26.70 7.43 0.81
FUDGE (Yang & Klein, 2021) can be viewed as the AR analog to our D-CBG. DiGress (Vignac et al., 2022) uses a similar first-order approximation to the one we employ. However, in DiGress this approximation is used to resolve the intractability of the normalizing constant. In our work, we derive a tractable expression for classifier-based guidance and simply use the Taylor approximation to speed up computation in large sequence length and vocabulary size regimes. Sanchez et al. (2023) derives an equivalent formulation for D-CFG and use it to better enforce AR models’ adherence to prefix prompts. FreeGress (Ninniri et al., 2024) also offer a comparable method to D-CFG, but focus on graph diffusion models. Most similar to our method, is the concurrent work of Nisonoff et al. (2024), which derives classifier-based and classifier-free guidance for discrete diffusion and flow models. However, this work is highly tailored to models that leverage the formalism of CTMC, and guidance is applied to the rate matrices.
@@ Appendix A Continuous Time Discrete Uniform Diffusion
Appendix B Relating UDLM to CTMC
Similar to Sahoo et al. (2024a), our work tackles the problem of discrete diffusion from the variational perspective and analyzes the ELBO in the continuous time limit. In contrast, other works, such as Campbell et al. (2022) and Lou et al. (2023), have extended the discrete diffusion framework proposed in Austin et al. (2021) using the formalisms of continuous time Markov chains (CTMC). In this section, we relate these two approaches.
@@ Appendix B Relating UDLM to CTMC
In the CTMC formulation, the key quantity of interest is the rate matrix $R_{t}\in\mathbb{R}^{N\times N}$. In analogy to the discrete time transition matrices $Q_{t}$, which define the transition probabilities between states, these rate matrices define the instantaneous rate of change between states in continuous time. More formally, letting $\delta$ be the Kronecker delta function, and using ${\mathbf{z}}$ and ${\mathbf{z}}^{\prime}$ to denote observed values of the latents, we can define the forward noising process using $R_{t}$ as follows:
$q({\mathbf{z}}_{t}={\mathbf{z}}^{\prime}\mid{\mathbf{z}}_{s}={\mathbf{z}})=\delta_{{\mathbf{z}}^{\prime},{\mathbf{z}}}+R_{t}({\mathbf{z}},{\mathbf{z}}^{\prime})\frac{1}{T}+o\Big{(}\frac{1}{T}\Big{)}$ (37)
where $o(1/T)$ indicates terms that vanish more quickly than $1/T.$ Recall that we defined $s=t-\frac{1}{T}.$ In continuous time, as $T\rightarrow\infty$, $o(1/T)$ terms are ignored, and we have
$R_{t}({\mathbf{z}},{\mathbf{z}}^{\prime})=\lim_{T\rightarrow\infty}\frac{q({\mathbf{z}}_{t}={\mathbf{z}}^{\prime}\mid{\mathbf{z}}_{t-(1/T)}={\mathbf{z}})-\delta_{{\mathbf{z}}^{\prime},{\mathbf{z}}}}{1/T},$ (38)
@@
Importantly, processes defined as in (37) have known time reversal processes given by (Kelly, 2011; Sun et al., 2022):
$q({\mathbf{z}}_{s}={\mathbf{z}}^{\prime}\mid{\mathbf{z}}_{t}={\mathbf{z}})=\delta_{{\mathbf{z}}^{\prime},{\mathbf{z}}}+\check{R}_{t}({\mathbf{z}},{\mathbf{z}}^{\prime})\frac{1}{T}+o\Big{(}\frac{1}{T}\Big{)},$ (39)
@@
$\check{R}_{t}({\mathbf{z}}^{\prime},{\mathbf{z}})=\frac{q({\mathbf{z}}_{t}={\mathbf{z}}^{\prime})}{q({\mathbf{z}}_{t}={\mathbf{z}})}R_{t}({\mathbf{z}},{\mathbf{z}}^{\prime}),~{}~{}~{}~{}~{}\text{if }{\mathbf{z}}^{\prime}\neq{\mathbf{z}}$ (40)

### v3

@@ 5.2 Guided Discrete Diffusion
AR FUDGE 924.4±${}_{\pm\text{12.1}}$ ${}_{\pm\text{3.5}}$ 0.61±${}_{\pm\text{0.00}}$ 281.2±${}_{\pm\text{5.0}}$ 103.6±${}_{\pm\text{4.8}}$ 4.89±${}_{\pm\text{0.10}}$
AR PPLM 1007.2±${}_{\pm\text{2.3}}$ ${}_{\pm\text{14.2}}$ 0.45±${}_{\pm\text{0.00}}$ 1009.8±${}_{\pm\text{1.9}}$ 140.6±${}_{\pm\text{16.5}}$ 1.92±${}_{\pm\text{0.09}}$
MDLM D-CBG 417.6±${}_{\pm\text{19.7}}$ 116.6±${}_{\pm\text{8.9}}$ 0.58±${}_{\pm\text{0.00}}$ 113.0±${}_{\pm\text{8.8}}$ 85.6±${}_{\pm\text{8.8}}$ 4.75±${}_{\pm\text{0.23}}$
MDLM NOS 506.0±${}_{\pm\text{29.5}}$ 240.4±${}_{\pm\text{16.0}}$ 0.45±${}_{\pm\text{0.01}}$ 247.8±${}_{\pm\text{14.0}}$ 193.6±${}_{\pm\text{13.1}}$ 3.51±${}_{\pm\text{0.22}}$
UDLM D-CBG 994.8±${}_{\pm\text{2.9}}$ 63.8±${}_{\pm\text{8.1}}$ 0.61±${}_{\pm\text{0.00}}$ 897.2±${}_{\pm\text{11.6}}$ 432.0±${}_{\pm\text{19.1}}$ 4.84±${}_{\pm\text{0.02}}$
UDLM NOS 547.8±${}_{\pm\text{10.6}}$ 158.8±${}_{\pm\text{11.0}}$ 0.47±${}_{\pm\text{0.00}}$ 573.8±${}_{\pm\text{13.0}}$ 244.4±${}_{\pm\text{13.1}}$ 3.96±${}_{\pm\text{0.07}}$
Table 6: Guidance improves FID & IS on CIFAR10. Finite- (D3PM) vs. continuous-time (MDLM / UDLM). Best values are bolded. From Austin et al. (2021).
FID ($\downarrow$) IS ($\uparrow$)
D3PM Absorb 41.28 6.26
MDLM 33.75 6.74
MDLM D-CFG 15.56 9.02
D3PM Uniform 51.27 5.99
UDLM 33.65 6.86
UDLM D-CFG 23.21 8.66
Molecular Property Maximization  For QM9, we investigate novel generation of sequences that maximize either drug-likeness (QED; Bickerton et al. (2012)) or number of rings present in the molecule. We only report values for which we generated at least 50 novel sequences (out of 1,024). In Table 5, for D-CFG, we see that all methods perform comparably when maximizing drug likeness, but that MDLM and UDLM are better suited for the structural property of ring count. For classifier-based guidance mechanisms, we again find that both QED and ring count, D-CBG with discrete diffusion best trades-off maximizing these properties while generating valid and novel sequences compared to AR or diffusion with NOS.
Figure 3: Diffusion models extend the steer-ability Pareto frontier. (Left) D-CBG outperforms FUDGE classifier guidance when maximizing drug-likeness (QED). (Right) D-CFG with diffusion better trades-off novel generation and ring-count maximization compared to AR.
@@
Ablation: Faster Sampling  In Table 7, we explore using faster inference settings (smaller $T$) where diffusion models predict multiple pixels in parallel. These results focus on the CIFAR10 dataset; see Appendix E.3 for additional experiments on other datasets. MDLM’s performance deteriorates with smaller $T$, whereas UDLM is robust to this setting. This validates a key motivation behind UDLM: in settings where MDLM ‘locks in’ certain predictions that it cannot change, UDLM is more resilient given that all tokens can change throughout the decoding process.
@@ 6 Related Works, Discussion, and Conclusion
Table 7: UDLM is robust to faster sampling. CIFAR10 images sampled from a conditional model (D-CFGγ=1). Best values per $T$ are bolded.
Model FID ($\downarrow$) IS ($\uparrow$)
@@
MDLM 27.94 7.14
UDLM 26.70 7.43
Discrete Diffusion  Recent works have examined interpolating discrete diffusion (Ou et al., 2024; Sahoo et al., 2024a; Shi et al., 2024; Zhao et al., 2024; Zheng et al., 2023), a special case of the general framework from D3PM (Austin et al., 2021). Our UDLM method most closely aligns to the state-of-the-art discrete diffusion of Ou et al. (2024), Sahoo et al. (2024a), and Shi et al. (2024) that focus on absorbing state diffusion. Similar to these works, we provide a continuous time NELBO leading to performance gains, but we focus on uniform noise. Of note, other extensions of D3PM that do not start from the variational perspective, instead rely on the formalisms of continuous time Markov chains (CTMC) (Campbell et al., 2022) and concrete score matching (Lou et al., 2023), but are less performant than works such as MDLM and our method. Also stemming from CTMC are extensions of flow-based (Lipman et al., 2022) approaches to discrete data (Campbell et al., 2024; Gat et al., 2024). In Appendix C, we discuss connections of our work to CTMC in more detail.
Guidance  Leveraging continuous embeddings of discrete data, Diffusion-LM (Li et al., 2022) uses Langevin sampling with CBG. Similarly, SSD-LM (Han et al., 2022) perform Gaussian noising on the logits of a bi-directional model, which they combine with pre-trained classifiers to perform CBG with Langevin dynamics. LD4LG (Lovelace et al., 2024) implement CFG on continuous embeddings. Stark et al. (2024) use flow-matching on the simplex to perform CFG and CBG on continuous representations. Wang et al. (2023) perform guidance using auxiliary semantic latent variables. In contrast, to these continuous formulations for discrete data, our work adapts guidance mechanisms directly to the discrete domain.
FUDGE (Yang & Klein, 2021) can be viewed as the AR analog to our D-CBG. DiGress (Vignac et al., 2022) uses a similar first-order approximation to resolve the intractability of the normalizing constant. In our work, we derive a tractable expression for classifier-based guidance and simply use the Taylor approximation to speed up computation in large sequence length and vocabulary size regimes. Sanchez et al. (2023) derives an equivalent formulation for D-CFG and use it to better enforce AR models’ adherence to prefix prompts. FreeGress (Ninniri et al., 2024) also offer a comparable method to D-CFG, but focus on graph diffusion models. Most similar to our method, is the concurrent work of Nisonoff et al. (2024), which derives classifier-based and classifier-free guidance for discrete diffusion and flow models. However, this work is highly tailored to models that leverage the formalism of CTMC, and guidance is applied to the rate matrices. See Appendix B for details on relating our guidance methods to Nisonoff et al. (2024).
@@ Appendix A Continuous Time Discrete Uniform Diffusion
Appendix B Relating Guidance to CTMC
Below we demonstrate how our proposed guidance mechanisms from Section 3 admit straightforward continuous-time extensions. In concurrent work, Nisonoff et al. (2024) use the formalisms of continuous time Markov chains (CTMC) to derive similar guidance formulas, which we show are equivalent to the continuous-time extensions of our work.
@@ Appendix B Relating Guidance to CTMC
In the CTMC formulation, the key quantity of interest is the rate matrix $R_{t}\in\mathbb{R}^{N\times N}$. In analogy to the discrete time transition matrices $Q_{t}$, which define the transition probabilities between states, these rate matrices define the instantaneous rate of change between states in continuous time. More formally, letting $\delta({\mathbf{x}},{\mathbf{y}})$ be the Kronecker delta function that equals 1 if ${\mathbf{x}}={\mathbf{y}}$ and 0 otherwise, and using ${\mathbf{z}}$ and ${\mathbf{z}}^{\prime}$ to denote observed values of the latents, we can define the forward noising process using $R_{t}$ as follows:
$q({\mathbf{z}}_{t}={\mathbf{z}}^{\prime}\mid{\mathbf{z}}_{s}={\mathbf{z}})=\delta_{{\mathbf{z}},{\mathbf{z}}^{\prime}}+R_{t}({\mathbf{z}},{\mathbf{z}}^{\prime})\frac{1}{T}+o\Big{(}\frac{1}{T}\Big{)}$ (37)
where $o(1/T)$ indicates terms that vanish more quickly than $1/T.$ Note that when clear from the context, we use the shorthand $\delta_{{\mathbf{z}},{\mathbf{z}}^{\prime}}:=\delta({\mathbf{z}},{\mathbf{z}}^{\prime}).$ Recall that we defined $s=t-\frac{1}{T}.$ In continuous time, as $T\rightarrow\infty$, $o(1/T)$ terms are ignored, and we have
$R_{t}({\mathbf{z}},{\mathbf{z}}^{\prime})=\lim_{T\rightarrow\infty}\frac{q({\mathbf{z}}_{t}={\mathbf{z}}^{\prime}\mid{\mathbf{z}}_{t-(1/T)}={\mathbf{z}})-\delta_{{\mathbf{z}},{\mathbf{z}}^{\prime}}}{1/T},$ (38)
@@
Importantly, processes defined as in (37) have known time reversals given by (Kelly, 2011; Sun et al., 2022):
$q({\mathbf{z}}_{s}={\mathbf{z}}\mid{\mathbf{z}}_{t}={\mathbf{z}}^{\prime})=\delta_{{\mathbf{z}}^{\prime},{\mathbf{z}}}+\check{R}_{t}({\mathbf{z}}^{\prime},{\mathbf{z}})\frac{1}{T}+o\Big{(}\frac{1}{T}\Big{)},$ (39)
@@
$\check{R}_{t}({\mathbf{z}}^{\prime},{\mathbf{z}})=\frac{q({\mathbf{z}}_{t}={\mathbf{z}}^{\prime})}{q({\mathbf{z}}_{t}={\mathbf{z}})}R_{t}({\mathbf{z}},{\mathbf{z}}^{\prime}),~{}~{}~{}~{}~{}\text{if }{\mathbf{z}}^{\prime}\neq{\mathbf{z}},$ (40)
@@
We remark that in the derivations for guidance in Nisonoff et al. (2024), they slightly change the form of (39) to:
$q({\mathbf{z}}_{s}={\mathbf{z}}\mid{\mathbf{z}}_{t}={\mathbf{z}}^{\prime})=\delta_{{\mathbf{z}}^{\prime},{\mathbf{z}}}(1+\check{R}_{t}({\mathbf{z}}^{\prime},{\mathbf{z}}^{\prime}))(1/T)+(1-\delta_{{\mathbf{z}}^{\prime},{\mathbf{z}}})\check{R}_{t}({\mathbf{z}}^{\prime},{\mathbf{z}})(1/T)+o(1/T).$ (41)
Sampling with Learned Reverse Rate Matrices
Assuming we have a learned model $R_{t,\theta}({\mathbf{z}}^{\prime},{\mathbf{z}})$ that approximates the true reverse rates (see Campbell et al. (2022) or Campbell et al. (2024) for details), we can generate from this model by drawing samples from the limiting distribution ${\mathbf{z}}_{1}\sim\bm{\pi}$ and using a forward Euler discretization to sample a chain of latents that end in samples ${\mathbf{x}}\sim q({\mathbf{x}})$ (Campbell et al., 2024):
${\mathbf{z}}_{s}\sim\mathrm{Cat}({\mathbf{z}}_{s};\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})+(1/T)\cdot R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})),$ (42)
where $1/T$ represents the discretization step size and, as above, $s=t-1/T.$ Additionally, note that here we follow the convention of diffusion models and let the time variable go from $t=1\rightarrow 0$ to indicate moving from noise towards signal, in contrast with the flow matching literature, where the reverse time convention is used (Lipman et al., 2022; Campbell et al., 2024).
B.1 CTMC Formulation of D-CFG
In Nisonoff et al. (2024), guidance is achieved by scaling the parametrized reverse rate matrix $R_{t,\theta}$. Predictor-free guidance (the analogous method to D-CFG) is presented as:
$R^{\gamma}_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}\mid y)=R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}\mid y)^{\gamma}\cdot R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})^{(1-\gamma)},$ (43)
where $R_{t,\theta}(\cdot\mid y)$ is the learned reverse rate when conditioned on $y$.
We now present the continuous-time extension of D-CFG and show that it is proportional to using the predictor-free guided reverse rate from (43) in conjunction with the forward Euler sampling from (42). Starting from the formulation of D-CFG from our work (where we make more explicit which variables are free / set)
$p^{\gamma}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t}={\mathbf{z}}^{\prime},y)\propto p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t}={\mathbf{z}}^{\prime},y)^{\gamma}\cdot p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t}={\mathbf{z}}^{\prime})^{(1-\gamma)},$
when using the forward Euler sampling scheme, we can plug in the formulation from (41) (with the true reverse rate replaced with the approximate one) into the right-hand side of this equation to get
$p^{\gamma}$ $({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t}={\mathbf{z}}^{\prime},y)\propto$
$\Bigg{[}\Big{(}\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})(1+R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}={\mathbf{z}}^{\prime}\mid y))(1/T)+(1-\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}))R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}\mid y)(1/T)\Big{)}^{\gamma}$
$+\Big{(}\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})(1+R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}={\mathbf{z}}^{\prime}))(1/T)+(1-\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}))R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})(1/T)\Big{)}^{(1-\gamma)}\Bigg{]}$
$=\Bigg{[}\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})(1+R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}={\mathbf{z}}^{\prime}\mid y))^{\gamma}\cdot(1+R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}={\mathbf{z}}^{\prime}))^{(1-\gamma)}(1/T)$
$~{}~{}~{}~{}~{}~{}+(1-\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}))R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}\mid y)^{\gamma}\cdot R_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})^{(1-\gamma)}(1/T)\Bigg{]}$
$=\Bigg{[}\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})(1+R^{\gamma}_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}={\mathbf{z}}^{\prime}\mid y))(1/T)$
$~{}~{}~{}~{}~{}~{}+(1-\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}))R^{\gamma}_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}\mid y)(1/T)\Bigg{]},$
where the first equality comes from the fact that all cross terms involving $\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}})\cdot(1-\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}))$ evaluate to zero, allowing us to aggregate terms under the exponents $\gamma$ and $1-\gamma,$ and the second equality comes from (43).
B.2 CTMC Formulation of D-CBG
While for classifier-free guidance, we did not need to distinguish between latent variables with a single token and those representing sequences of tokens (as argued in Section 3.1), below we make this distinction explicit by including superscripts $(\ell)$ and $(1:L)$ on variables ${\mathbf{z}}_{s},{\mathbf{z}}_{t}.$ However, for notational simplicity we omit theses subscripts on variables that denote realized values, such as ${\mathbf{z}}^{\prime}$.
In Nisonoff et al. (2024), predictor guided rates (the analog to our D-CBG) are attained by scaling $R_{t,\theta}$ as follows:
$R_{t,\theta}^{\gamma}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)}\mid y)=\Bigg{(}\frac{p_{\phi}(y\mid{\mathbf{z}}_{s}^{(1:L)},{\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime})}{p_{\phi}(y\mid{\mathbf{z}}_{s}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime})}\Bigg{)}^{\gamma}R_{t,\theta}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)}),$ (44)
where $p_{\phi}$ is an external classifier. Given the factorization assumptions of the forward and reverse processes, the reverse rate matrix in the CTMC formulation is only non-zero at entries where its arguments differ in (at most) one dimension (Campbell et al., 2022; 2024; Nisonoff et al., 2024). Thus, the classifier only needs to be evaluated on candidates for which latents differ in at most one token, which coincides with the argument / classifier we introduce in Section 3.2. The formulation in (44) can be further simplified to (see Nisonoff et al. (2024) for details):
$R_{t,\theta}^{\gamma}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)}\mid y)=\frac{p_{\phi}(y\mid{\mathbf{z}}_{s}^{(1:L)})}{p_{\phi}(y\mid{\mathbf{z}}_{s}^{(1:L)}={\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime})}R_{t,\theta}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)}),$ (45)
where for each $\ell$, using the notation introduced in Section 3.2, we only evaluate $p_{\phi}$ on ${\mathbf{z}}_{s}^{(1:L)}\in\tilde{\mathcal{Z}}_{\ell}({\mathbf{z}}_{t}^{(1:L)}).$
As in Appendix B.1, we present the continuous-time extension of D-CBG and show that it is proportional to sampling using the reverse rate matrix in (45) (with Euler discretization for sampling).
Let $R_{t,\theta}^{(\ell)}$ represent the rate matrix corresponding to transitions where all dimensions are kept the same except for potentially the one corresponding to the $\ell^{\text{th}}$ token. We follow a similar derivation to that used in Appendix B.1 and start with our formulation for D-CBG:
$p^{\gamma}({\mathbf{z}}_{s}^{(\ell)}\mid{\mathbf{z}}_{t}^{(1:L)},y)\propto p_{\phi}(y\mid\tilde{{\mathbf{z}}}^{(1:L)})^{\gamma}p_{\theta}({\mathbf{z}}_{s}^{(\ell)}\mid{\mathbf{z}}_{t}^{(1:L)}),$
where recall that $\tilde{{\mathbf{z}}}^{(1:L)}=\Big{[}{\mathbf{z}}_{t}^{(1:\ell-1)},{\mathbf{z}}_{s}^{(\ell)},{\mathbf{z}}_{t}^{(\ell+1:L)}\Big{]}$. We then plug in the formulation for the approximate reverse rate to get:
$p^{\gamma}({\mathbf{z}}_{s}^{(\ell)}\mid{\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},y)\propto$
$~{}p_{\phi}(y\mid\tilde{{\mathbf{z}}}^{(1:L)})^{\gamma}\cdot\Bigg{(}\delta({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)})(1+R^{(\ell)}_{t,\theta}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)}={\mathbf{z}}^{\prime}))(1/T)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+(1-\delta({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)}))R^{(\ell)}_{t,\theta}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)})(1/T)\Bigg{)}$
$\propto\Bigg{(}\frac{p_{\phi}(y\mid\tilde{{\mathbf{z}}}^{(1:L)})}{p_{\phi}(y\mid{\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime})}\Bigg{)}^{\gamma}\cdot\Bigg{(}\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s})(1+R^{(\ell)}_{t,\theta}({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}={\mathbf{z}}^{\prime}))(1/T)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+(1-\delta({\mathbf{z}}_{t}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}))R^{(\ell)}_{t,\theta}({\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime},{\mathbf{z}}_{s}^{(1:L)})(1/T)\Bigg{)},$
where the second proportionality comes from the fact that $p_{\phi}(y\mid{\mathbf{z}}_{t}^{(1:L)}={\mathbf{z}}^{\prime})$ is simply a constant. Therefore, we have shown the equivalence between our D-CBG method and the predictor guidance method introduced in Nisonoff et al. (2024).
Appendix C Relating UDLM to CTMC
Similar to Sahoo et al. (2024a), our work tackles the problem of discrete diffusion from the variational perspective and analyzes the ELBO in the continuous time limit. In contrast, other works, such as Campbell et al. (2022) and Lou et al. (2023), have extended the discrete diffusion framework proposed in Austin et al. (2021) using the formalisms of CTMC. In this section, we relate these two approaches.